OpenStudy (anonymous):

can someone please explain to me why the answer to the following question is -infinity, infinity

OpenStudy (anonymous):

ticktock

OpenStudy (anonymous):

\[\lim_{x \rightarrow -infinity} (\ln(cbrt(x)))/sinx\]

OpenStudy (anonymous):

cbrt = cube root

OpenStudy (anonymous):

lnx^(1/3)/sinx = (1/3)lnx / sinx ? i thinkk..

OpenStudy (anonymous):

actually apparently -infinity, infinity is not the correct answer.... wolfram lied so if you help me come to the correct answer

OpenStudy (anonymous):

?? \[\lim_{x\to -\infty}\frac{\ln(\sqrt[3]{x})}{\sin(x)}\]

OpenStudy (anonymous):

ya thats it

OpenStudy (anonymous):

makes no sense since you cannot take the log of a negative number

OpenStudy (anonymous):

i think should be 0

OpenStudy (anonymous):

so there is no limit, as the numerator is undefined if \(x\leq 0\)

OpenStudy (anonymous):

oh crap it is \[\lim_{x \rightarrow +infinity}\] sorry....

OpenStudy (anonymous):

\(\sin(x)\) has no limit either as it is periodic and takes on all values between -1 and 1 infinitely often

OpenStudy (anonymous):

well still no limit

OpenStudy (anonymous):

numerator is going to infinity, but denominator is not going to any specific number as i wrote above.

OpenStudy (anonymous):

in other words it swings wildly between large positive and large negative values as sine varies between -1 and 1

OpenStudy (anonymous):

so ultimately its undefined

OpenStudy (anonymous):

yes for sure undefined

OpenStudy (anonymous):

so with that being said, anytime sinx is in my denomonator i know that the function must be undefined?

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

no

OpenStudy (anonymous):

well no, the numerator could go to zero, so the whole thing could go to zero

OpenStudy (anonymous):

\[\lim_{x\to\infty}\frac{e^{-x}}{\sin(x)}\] for example would be zero

OpenStudy (anonymous):

so are there any good tips and tricks you could tell me to help me understand this sort of material

OpenStudy (anonymous):

no tricks really, you have to make sure you know what it going on, and don't for example us l'hopital's rule unless it is applicable, i.e. unless you are in indeterminate form, which you are not here. in your example you do not have \(\frac{\infty}{\infty}\) and so you cannot use l\hopital and get zero for an answer!

OpenStudy (anonymous):

see im not sure how to determine whether i am in that form or not

OpenStudy (anonymous):

plug in the number and check or imagine what happens as x gets large if you are going to infinity

OpenStudy (anonymous):

in your example as x gets large so does \(\sqrt[3]{x}\) and therefore so does \(\ln(\sqrt[3]{x})\) but the prolem is that sine does not get large and does not have a limit

OpenStudy (anonymous):

so basically what i am saying is that in this problem there is no trick or gimmick. you just have to imagine what happens as x gets bigger and bigger

OpenStudy (anonymous):

ok thanks!